# Optimizing Multi-level Magic State Factories for Fault-Tolerant Quantum Architectures

### v1

@@ III Hardware Error Modelling
Designing the fault-tolerant quantum architecture described requires solving a bi-objective optimization problem that balances minimizing the allocated space (i.e., the physical qubits required) and time (i.e., the runtime for circuit execution) under a given error budget. The decisions to be made are related to sizing the components of the architecture—specifically, the core processor and the MSF—in such a way that both objectives are optimized while ensuring fault tolerance.
@@
In this section we show how the errors of the code processor and the MSF are modelled and predicted. Although generalizing our models and methods for more-complex architectures is a straightforward task, we present them considering that similar distillation units (e.g., layout and protocol) are used across all levels of the MSF for simplicity. A table summarizing the definition of all variables and parameters used by our error models is given as Appendix A.
III.1 Core Processor Errors
In the core processor, the quantum algorithm operates on $Q$ logical qubits to execute $T$ non-Clifford gates (i.e., $\pi/8$ Pauli rotations) over a makespan of $K$ logical cycles. The cycles are divided into active cycles, where at least one non-Clifford gate is performed in the memory fabric as determined by a compilation procedure, and idling cycles, where the data qubits in the memory fabric must be preserved through error correction while waiting for resources (e.g., magic states).
During active cycles, lattice surgeries connect magic states from the buffer to the data qubits to perform the auto-corrected non-Clifford gates scheduled for that cycle. Lattice surgeries used to send magic states from the growth units to the buffers are also considered, although they are performed in previous cycles, either active or idling ones. Each of these lattice surgeries require performing a logical operation involving faulty logical qubits. The error probability for all lattice surgeries required to perform an operation with index $i\in\{1,\ldots,T\}$ is denoted as $e_{\text{surg},i}$, and depends on the size and structure of the lattice surgeries performed [18].
$V_{\text{idle}}=KQ-\sum\limits_{i=1}^{T}q_{i},$ (2)
where $KQ$ represents all locations where a logical error can occur when protecting idling qubits, and $q_{i}$ is the number of logical data qubits involved in each non-Clifford gate, which is subtracted from the total to avoid double-counting of the data qubits that are actively involved in a lattice surgery.
$E_{\text{core}}=V_{\text{idle}}e_{\text{mem},\text{core}}+\sum\limits_{i=1}^{T}e_{\text{surg},i}.$ (3)
The above formula can be approximated as
$E_{\text{core}}=V_{\text{idle}}e_{\text{mem},\text{core}}+Te_{\text{cliff},\text{core}}.$ (4)
where $\alpha$ is the average number of patches in all lattice surgeries in the compiled program and $e_{\text{cliff},\text{core}}$ represents the average error rates of all the multi-qubit lattice surgeries in the core. This approximation disregards the varying shapes of different lattice surgeries.
An alternative further simplification is to only use quantum memory error rates. To this end we define the active volume of a compiled program as
$V_{\text{act}}=(2Q+\sqrt{8Q}+26)\alpha T,$ (5)
where $(2Q+\sqrt{8Q}+26)$ refers to the number of logical qubits in the core, and $\alpha$ is the average size (i.e., number of surface code patches) of the lattice surgeries in the program. This results in the approximation
We assume that the assembler has access to the logical error rates of the QEC codes (at high distances). This information may be provided directly by experiments on the quantum hardware. Alternatively, a predictive (parameterized) model $f_{\text{mem}}$ can be regressed from numerical simulations at low distances using efficient stabilizer circuit simulators [19, 20]. To summarize, for a core processor of distance $d_{\text{core}}$ we have
$e_{\text{mem},\text{core}}=f_{\text{mem}}(d_{\text{core}}).$ (7)
as our model, where the error prefactor $\mu_{\text{mem}}$ and the error suppression rate $\Lambda_{\text{mem}}$ are fitting parameters. The error suppression rate $\Lambda_{\text{mem}}$ represents the reduction in the logical error rate when increasing the code distance by two. We expect $\Lambda_{\text{mem}}>1$ for the surface code below threshold [18, 21, 8]. The exponential suppression with base $\Lambda$ represents the length $(d+1)/2$ of the most probable error strings, and the coefficient $d^{2}$ accounts for the multiplicity of these error strings, governed by the area of the cross section of the code bulk with error strings of either $X$ or $Z$ type as shown in Fig. 4(a).
(a) Most probable quantum memory errors
(b) Most probable X-type errors in surgery
(c) Most probable Z-type errors in surgery
(d) Most probable time-like errors in surgery
Figure 4: Shortest-length undetectable errors in FTQC protocols determine the leading factors of the predictive error models. (a) For quantum memory experiments both $X$ and $Z$ type errors land on boundaries of area $rd$ for $r$ rounds of stabilizer measurements. (b-d) For mult-qubit surgeries, no extra rounds of QEC is assumed at the split and the pre-merge phases. (b) Shortest-length error strings of the $X$-type have a cross section of area $O((2d+b)r)$. (c) Since no distinct pre-merge rounds are assumed, the area of the boundaries of the shortest-length $Z$-type errors grows as $O(d)$. (d) The degeneracy of the shortest-length undetectable time-like errors scales with the area of the measured bus, $db$.
Similarly, the Clifford error rates $e_{\text{cliff}}$ can be approximated using stabilizer circuit simulations at low distances and extrapolated to higher distances. A model for the multi-qubit surgery errors can be extracted from a similar argument involving the shortest length errors in the bulk as shown in Fig. 4. Note that an area law does not apply to all error string types [22]. For example, in an $XX$ surgery between two logical qubits involving $r$ QEC rounds and a bus patch of length $b$, the error model is
since the $X$-type errors have boundary surfaces of areas $r(2d+b)$ (see Fig. 4(b)), the $Z$-type errors grow with the length of the boundary edges shown in Fig. 4(c) and therefore contribute to a linear factor of $d$ in the second term, and the time-like errors of the shortest distance land in a boundary of area $bd$ (Fig. 4(d)).
For teleporations in our core processor, the teleported state is a magic state that has resided in the buffer for some average expected buffer delay time, $\tau_{b}$ (which is 1 clock cycle for balanced production and consumption rates; see Section IV.1 for further details). The targets of teleporation are logical data qubit patches in the core processor for which further QEC rounds are executed until the decoder outcome is available. We denote this delay by $\tau_{d}$. Inclusion of the buffer and decoder delays and assuming an average rate for all types of surgeries changes the above model to
which still distinguishes time-like and space-like errors but ignores the type of surgery; e.g., $XX$ merge or otherwise (Fig. 5).
Figure 5: The effect of buffer and decoder delays on surgery fidelities. In core processor surgeries, a teleportation is implemented involving a magic state (the left code patch) and data qubits (represented by the right code patch). The magic state incurs a delay $\tau_{b}$ in the buffer, which in the space-efficient steady flow scenario is one logical cycle on average (i.e., $d$ QEC rounds) before the surgery starts. The data qubit QEC rounds must continue until decoder decisions are available after decoder delay time $\tau_{d}$.
III.2 Magic State Factory Errors
$E_{\text{msf}}=Te_{\text{msf}},$ (11)
where $e_{\text{msf}}$ is the error rate of the magic state consumed by each operation in the core.
In the MSF, an iterative process is followed that involves multiple levels of distillation to improve fidelity linked with each other using growth zones. Consider an MSF with $L$ levels: a 00-th level referring to the magic state preparation area, the subset $\{1,\ldots,L\}$ referring to the consecutive distillation levels, and the $L+1$-st level being to the core processor. The number of distillation levels $L\in\mathbb{Z}_{\geq 0}$ is to be determined such that the magic state error rate $e_{\text{msf}}$ meets the condition in Eq. 1 while satisfying Eq. 11. And the number of qubits used for building the MSF (the space cost) is optimized by choosing an increasing sequence $d_{0},d_{1},\ldots,d_{L+1}=d_{\text{core}}$ of code distances for each of the levels assuming a fault-tolerant code growth protocol is applied within the growth zones to promote code patches of distance $d_{\ell}$ to those of distance $d_{\ell+1}$.
The logical error rate of the magic states provided to the $\ell$-th MSF level is denoted by $e_{\text{in},\ell}$ and the error rate of the resulting distilled magic state is $e_{\text{out},\ell}$. Therefore the contribution of the growth zones to the error rates of the magic states between levels is calculated via
Similar to Section III.1 a predictive model (again, to be determined theoretically and regressed numerically or alternatively experimentally) simulating the specific growth protocol is used to infer
$e_{\text{grow},\ell}=f_{\text{grow}}(d_{\ell},d_{\ell+1})$ (13)
$e_{\text{out},\ell}=g_{\text{dist}}(e_{\text{in},\ell},e_{\text{cliff},\ell})$ (14)
which for example for the 15:1 distillation protocol is
$e_{\text{out},\ell}=35e_{\text{in},\ell}^{3}+7.1e_{\text{cliff},\ell}$ (15)
as per Refs. [23, 5, 11]. The coefficient 35 is driven theoretically [23] however the contribution of surgeries within the $\ell$-th distillation circuit is approximated as a numerically driven multiple (7.1 in [24]) of the Clifford error rate $e_{\text{cliff},\ell}$ obtained from the model Eq. 10.
$e_{\text{out},0}$ $=e_{\text{in},1}=e_{\text{prep}},$ (16)
Here the preparation error rate $e_{\text{prep}}$ is estimated from numerical simulation of the fault-tolerant magic state preparation protocol in use at the zeroeth MSF level (i.e., at distance $d_{0}$). The numerical behaviour of these protocols as a function of $d$ in the error regimes in which we are interested can be quite distinct. For instance, for some protocols [25, 17], the error rate
with a negative error suppression rate $\Lambda_{\text{prep}}$, which leads to the error rate increasing linearly with respect to distance. However, more recent protocols [26] exhibit an exponential suppression in their error rates as far as numerical simulations in the regime of practically relevant distances and physical error rates are concerned:
$e_{\text{prep}}=\mu_{\text{prep}}\Lambda_{\text{prep}}^{-\frac{d+1}{2}}.$ (21)
Combining Eqs. 2, 3 and 11 we can rewrite the error budget constraint Eq. 1 as
$e_{\text{msf}}T+e_{\text{mem},L+1}\left(KQ-\sum\limits_{i=1}^{T}q_{i}\right)+\sum\limits_{i=1}^{T}e_{\text{surg},i}\leq E.$ (22)
Therefore, given a compiled quantum circuit with $T$ non-Clifford gates involve $Q$ logical qubits, run (potentially with some parallelization) in $K$ logical cycles, involving $q_{i}$ qubits in the $i$-th non-Clifford operation, and the hardware error models described above, we can determine the number of distillation levels $L$, the code distances at each level, and the code distance in the core processor such that Eq. 22 is satisfied.
IV The Assembly Process
In this section we show how the predicted error rates from the previous section are used to decide the sizing of the MSF and all code distances involved.
IV.1 Space and Time Costs
@@
The space and time costs with respect to the architecture depend on the sizes of the core processor and the MSF. Each logical cycle in the makespan $K$, either active or idling, requires performing $d_{L+1}$ parity checks, each taking a time $W$ determined by hardware characterization for the QEC scheme considered. Therefore, ignoring the warm-up time (i.e., the time needed to start filling the core buffer with high quality magic states), which is negligible for utility-scale quantum circuits, we can calculate the runtime $R$ of the quantum circuit as:
$R=KWd_{L+1}.$ (23)
Therefore, for a given QPU, $R$ is minimized by reducing the product of the code distance of the core processor $d_{L+1}$ and the makespan $K$, which is determined based on the production rate of magic states in the MSF and the demand profile for magic state consumption in the core.
@@
A compilation process, such as those described in Refs. [4] and [24], can determine sets of $\pi/8$ rotation gates that can be performed in parallel. The compilation in these studies assumes that magic states are always available to be consumed by all gates scheduled for the same logical cycle. Let a quantum algorithm composed of $T$ non-Clifford gates be represented by $T=\sum_{i=1}^{T_{\text{depth}}}t_{i}$, where $T_{\text{depth}}$ determines the minimum number of logical cycles required to execute the algorithm (the shortest possible makespan, $K\leq T_{\text{depth}}$), and $t_{i}$ denotes the maximum number of $\pi/8$ rotations that can be performed in the time step $i$ as determined by a compilation process. Ignoring fluctuations in the demand for magic states, we can define the maximum consumption rate of the core processor as
@@
since logical cycles would be expected to consume, on average, $T/T_{\text{depth}}$ magic states every logical cycle that takes a time $Wd_{L+1}$.
@@
$D_{\ell}\leq\dfrac{u_{\ell}NP}{Wd_{\ell}(O+N)},$ (26)
where $N$ is the number of magic states produced in a distillation round, $P$ is the acceptance probability, $Wd_{\ell}$ is the execution time of each logical operation in the unit, and $O$ is the number of logical cycles required for the distillation round. The parameters $N$ and $O$ are given based on the distillation protocol, the acceptance probability is a function $P=f_{\text{acc}}(e_{\text{in},\ell},e_{\text{mem},\ell})$, and $W$ is based on the hardware characterization. Equality in Eq. 26 is obtained when a steady execution of logical operations is possible due to a sufficient supply of magic states. Given that the consumption rate of magic states by all units in this level is
$C_{\ell}\leq\dfrac{u_{\ell}M}{Wd_{\ell}(O+N)},$ (27)
a steady flow is possible when $D_{\ell-1}\geq C_{\ell}$, where $M$ is the number of input magic states for the distillation unit. For example in the 15:1 distillation protocol $M=15$, $N=1$, $O=13$, and
@@
$\dfrac{u_{L}NP}{d_{L}(O+N)}$ $=\dfrac{d_{L+1}T_{\text{depth}}}{T},$ (30)
@@
Solving Eqs. 30 to 32 considering the error budget in Eq. 22 leads to an architecture that approximates the space required to minimize time costs. Space costs are measured by the number of physical qubits required to encode all logical qubits in the layout. The exact number of logical qubits in the core processor is determined as described in Section II.1 for the memory fabric, while we consider a buffer with a fixed size of 25 logical qubits (represented in Fig. 2) that guarantees a steady flow of magic states for a circuit with $T=T_{\text{depth}}$. And, each logical qubit in the surface code requires $12d^{2}-$ 1 physical qubits.
For example, in the 15:1 distillation unit layout used in Fig. 3:
@@
  • each distillation unit requires 23 logical qubits regardless of its level;
  • each distillation unit in all levels but the lowermost still requires 5 dedicated growth units supplying magic states to it, each requiring a space equivalent to a logical qubit with code distance $d_{\ell}$;
  • each distillation unit in the lowermost level requires 10 logical qubits for handling the prepared magic states, 3 of which are for preparation units as determined by preliminary experiments using common hardware characterizations; and
@@
  • each distillation level $\ell$ still requires at least $45u_{\ell}+$ 4 qubits to connect the distillation units to the upper-level growth units.
@@
Based on these considerations, the total space cost for a time-optimal architecture is approximated as
$S\approx(2d_{L+1}^{2}-1)(2Q+\sqrt{8Q}+29)+(2d_{1}^{2}-1)(41u_{1}+4)+\sum\limits_{\ell=2}^{L}(2d_{\ell}^{2}-1)(33u_{\ell}+4).$ (33)
@@
While solving Eqs. 30 to 32 leads to a time-optimal solution, if a distillation unit can meet the requirements to produce higher-fidelity magic states from lower-fidelity ones, only one distillation unit per level is necessary for an MSF to be able to produce magic states at the error rates meeting the condition in Eq. (22). As previously mentioned, an MSF with $u_{\ell}=1,\forall\ell\in\{1,\ldots,L\}$, may significantly reduce the last two terms in Eq. 33 at the cost of increasing the first one through an increase in $d_{L+1}$. This trade-off is analyzed in detail in Section V.
IV.2 Our Assembly Method
@@
Although the problem presented in Section III can be solved analytically, the complexity of representing the nested function for the accumulated MSF errors in Eq. 22 and the relations between the decision variables in Eqs. 30 to 32 demand the development of a fast heuristic algorithm to solve the problem. This algorithm generates layouts for the considered architecture under both time-optimal and -suboptimal scenarios.
@@
IV.2.1 Set the Core Area’s Code Distance

### v2

@@ III The Assembly Process
Designing the fault-tolerant described quantum architecture requires solving a bi-objective optimization problem that balances minimizing the allocated space (i.e., the physical qubits required) and time (i.e., the runtime for circuit execution) under a given error budget. The decisions to be made are related to sizing the components of the architecture—specifically, the core processor and the MSF—such that both objectives are optimized while ensuring fault tolerance.
@@
The error models detailed in Appendix B provide a breakdown of this error budget as
$e_{\text{msf}}T+e_{\text{mem},L+1}\left(KQ-\sum\limits_{i=1}^{T^{+}}q_{i}\right)+\sum\limits_{i=1}^{T^{+}}e_{\text{surg},i}\leq E.$ (2)
where $e_{msf}$ represents the error rate of each magic state consumed in the core processor, $e_{\text{mem},L+1}$ accounts for logical memory errors in the core processor, and $e_{\text{surg},i}$ denotes the error contribution from each lattice surgery operations in the core processor for a compilation with $K$ logical time steps (makespan) for all $T$ non-Clifford and $T^{+}-T$ Clifford operations in the circuit. The number of logical data qubits $q_{i}$ among all $Q$ data qubits involved in the logical quantum program is subtracted from the total to avoid double-counting the data qubits that are actively involved in a lattice surgery (see also Appendix D for a summary of all variables and parameters used in our error models). In this section, we show how the sizing of the MSF and all code distances involved are decided such that the accumulated errors from all quantum operations in the core processor and in the MSF are below $E$ while balancing space and time costs.
III.1 Space and Time Costs
@@
The space and time costs with respect to the architecture depend on the sizes of the core processor and the MSF, as well as the protocol used for the computation. Ignoring the warm-up time (i.e., the time needed to start filling the core buffer with high-quality magic states and prepare the correction qubits in the buffer), which is negligible for utility-scale quantum circuits, we can calculate the runtime $R$ of the quantum circuit as
$R=K\max{\{Wd_{L+1},\gamma\}}.$ (3)
In a $T$-count-limited computation, the makespan $K=T$ logical time steps, and each step will take the longer between a logical cycle time in the core processor, that is, $d_{L+1}$ parity checks, each taking a time $W$ determined by hardware characterization for the QEC scheme considered, and the reaction time $\gamma$. Idling logical cycles also require performing $d_{L+1}$ parity checks to protect the important data. In a reaction-time-limited computation, which is only optimal when $\gamma<Wd_{L+1}$ (see Section IV.4), $K=\delta\lceil T/(n-1)\rceil$ for an architecture with $n$ memory fabrics and each of the $\lceil T/(n-1)\rceil$ $BP$-$P$-$BM$ rounds taking $\delta$ logical time steps. For a given hardware device, $R$ is minimized by reducing the product of the code distance of the core processor $d_{L+1}$ and the makespan $K$, which is determined based on the production rate of magic states in the MSF and the demand profile for magic state consumption in the core.
Transpilation procedures can reduce circuit depth—and consequently the makespan—by commuting and eliminating Clifford gates from a quantum algorithm [4]. This optimization smooths fluctuations in magic state demand, enabling a more consistent consumption rate. Compilation methods, such as those described in Refs. [4] and [12], identify sets of $\pi/8$ rotation gates that can be performed in parallel. These approaches typically assume an uninterrupted supply of magic states for all gates scheduled within the same logical cycle.
@@
Let a quantum algorithm composed of $T$ non-Clifford gates be represented by $T=\sum_{i=1}^{T_{\text{depth}}}t_{i}$, where $T_{\text{depth}}$ denotes the minimum number of logical cycles required to execute the algorithm (the shortest possible makespan, $K\leq T_{\text{depth}}$), and $t_{i}$ is the maximum number of $\pi/8$ rotations executable in time step $i$, as determined by the compilation. By assuming a transpiled algorithm with only non-Clifford gates and a steady magic state supply system and ignoring demand fluctuations since transpiled circuits are generally less parallelizable, the core processor’s maximum consumption rate can be defined as
@@
since logical cycles would be expected to consume, on average, $T/T_{\text{depth}}$ magic states every logical cycle that takes a time $Wd_{L+1}$. For the reaction-time-limited computation, the consumption rate of each core processor is slowed down by a factor of the time required for the Bell state preparations and basis measurements. On average, assuming $T=T_{depth}$, $n-1$ magic states are consumed across all $n$ core processors every $\delta$ time steps.
@@
$D_{\ell}\leq\dfrac{u_{\ell}NP}{Wd_{\ell}O},$ (6)
where $N$ is the number of magic states produced in a distillation round, $P$ is the probability of success for magic state distillation, $Wd_{\ell}$ is the execution time of each logical operation in the unit, and $O$ is the number of logical cycles required for the distillation round. The parameters $N$ and $O$ are given based on the distillation protocol, the success probability is a function $P=f_{\text{acc}}(e_{\text{in},\ell},e_{\text{mem},\ell})$, and $W$ is based on the hardware characterization. Equality in Eq. 6 is obtained when a steady execution of logical operations is possible due to a sufficient supply of magic states. Given that the consumption rate of magic states by all units in this level is
$C_{\ell}\leq\dfrac{u_{\ell}M}{Wd_{\ell}O},$ (7)
a steady flow is possible when $D_{\ell-1}\geq C_{\ell}$, where $M$ is the number of input magic states for the distillation unit. For example in the 15:1 distillation protocol $M=15$, $N=1$, $O=13$ (i.e., 11 logical cycles for the $T$ gates in the protocol, one logical cycle for the corrections, and one cycle to empty the distilling port and position the lower-fidelity magic states for the next distillation cycle), and
@@
$\dfrac{u_{L}NP}{d_{L}O}$ $=\dfrac{d_{L+1}T_{\text{depth}}}{T},$ (10)
@@
Solving Eqs. 10 to 12 considering the error budget in Eq. 2 leads to an architecture that approximates the space required to minimize time costs. Space costs are measured by the number of physical qubits required to encode all logical qubits in each area of the layout. Let the total space cost $S$ be represented as the sum of the physical qubits required at each area $\ell$, i.e.,
$S=\sum\limits_{\ell=1}^{L+1}S_{\ell}.$ (13)
The number of logical qubits in the core processor is determined as described in Section II.1 for the memory fabric, that is, $292Q+\sqrt{8Q}+$ 29 logical qubits when $Q$ is a perfect square and slightly fewer logical qubits otherwise. We consider buffers with a size of 18 logical qubits (represented in Fig. 1), which guarantees a steady flow of magic states to the memory fabric at a rate of one magic state per logical cycle in the time-optimal case. The correction storage required is determined by converting the reaction time $\gamma$ into logical cycles and accounting for accessibility, approximated as one bus patch for every two additional correction storage patches. Therefore, given that each logical qubit in the surface code requires $12d^{2}-$ 1 physical qubits, the core processor requires approximately
$S_{L+1}=(2Q+\sqrt{8Q}+47+1.5\lceil\gamma/(d_{L+1}W)\rceil)(2d_{L+1}^{2}-1)$ (14)
physical qubits.
The total space cost for the MSF is given by the space required for each distillation level. In the 15:1 distillation unit layout used in Fig. 2:
  • each first-level distillation unit requires 29 logical qubits, plus additional qubits for magic state preparation and correction storage. Three preparation areas are used: the core requires a magic state prepared per time step, while the two upper areas require one every $O$ logical cycles. Using a steady-state approximation, the required number of preparation units per area is $n_{\text{prep}}=s\times(c_{\text{prep}}+p_{\text{prep}})/p_{\text{prep}}$, where $p_{\text{prep}}$ is the preparation unit success probability, $c_{\text{prep}}$ is the preparation cycle time, and $s$ the required production rate;
@@
  • each higher-level distillation unit requires 28 logical qubits, plus correction storage; and
@@
  • each distillation level $\ell$ still requires at least $110(u_{\ell}-1)+$ 1 qubits ($111u_{\ell}+$ 1 for the first level) to connect the distillation units to the upper-level growth units.
@@
Based on these considerations, the total space cost for the first level in the MSF is
$S_{1}=u_{1}(40+15/13(c_{\text{prep}}+p_{\text{prep}})/p_{\text{prep}}+1.5\lceil\gamma/(d_{L+1}W)\rceil)-10,$ (15)
and for each of the higher levels =$\ell=\{2,\ldots,L\}$, it is
$S_{\ell}=u_{\ell}(38+1.5\lceil\gamma/(d_{L+1}W)\rceil)-9.$ (16)
@@
While solving Eqs. 10 to 12 leads to a time-optimal solution, if a distillation unit can meet the requirements to produce higher-fidelity magic states from lower-fidelity ones, only one distillation unit per level is necessary for an MSF to be able to produce magic states at the error rates meeting the condition in Eq. (2). As previously mentioned, an MSF with $u_{\ell}=1,\forall\ell\in\{1,\ldots,L\}$, may significantly reduce the size of the MSF $\sum_{\ell=1}^{L}S_{\ell}$ at the cost of increasing $S_{L+1}$ through an increase in $d_{L+1}$. This trade-off is analyzed in detail in Section IV.
III.2 Our Assembly Method
@@
Although the problem presented in Appendix B can be solved analytically, the complexity of representing the nested function for the accumulated MSF errors in Eq. 2 and the relations between the decision variables in Eqs. 10 to 12 demand the development of a fast heuristic algorithm to solve the problem. This algorithm generates layouts for the considered architecture under both time-optimal and time-suboptimal scenarios.
@@
III.2.1 Set the Core Area’s Code Distance